OpenStudy (anonymous):

Construct a polynomial function f with the given properties: Zeros: -5(multiplicity 2), 2(multiplicity 1), 4(multiplicity 1); degree 4; contains point (3, 128). Im stuck on this halfway through..

OpenStudy (campbell_st):

y =(x+5)^2(x -2)(x-4) + c

OpenStudy (anonymous):

my book says -2(x+5)^2(x-2)(x-4)?

OpenStudy (anonymous):

i think i'm just confused on finding c.. how did it end up being -2?

OpenStudy (campbell_st):

the method is y = c(x+5)^2(x -2)(x -4) sub x = 3 and y = 128 then -64c = 128 c = -2

OpenStudy (campbell_st):

my mistake

OpenStudy (anonymous):

ohhhh wow I completely forgot about that.. I couldn't remember how to make sure the point was included but thats exactly what I've been looking for lol thanks!